OpenStudy (anonymous):

A dam released 1000 cubic meters of water at a rate of 10 cubic meters per minute, and then released another 1000 cubic meters at a rate of 20 cubic meters per minute. What was the avg rate at which the water was released? How do I go about this question ?

OpenStudy (anonymous):

1000/10 = 100 1000/20 = 50 100+50 = 150 2000/150 = 40/3 = 13.33 m^3/s